Find the vertex and focus of the parabola y^2-8x=0
Answer by lwsshak3(11628)

Find the vertex and focus of the parabola y^2-8x=0
**
This is a parabola that opens rightwards.
Its standard form of equation: (y-k)^2=4p(x-h), (h,k)=(x,y) coordinates of the vertex.
For given parabola:
y^2-8x=0
y^2=8x
vertex: (0,0)
axis of symmetry: x-axis or y=0
4p=8
p=2
focus: (2,0) (p distance from vertex on axis of symmetry)